# Sim2Real Transfer for Audio-Visual Navigation with Frequency-Adaptive Acoustic Field Prediction

### v1

@@ V Robot Platform
To deploy our sim2real policy on a real robot, we build our own audio-visual robot by equipping a HelloRobot with a 3Dio binaural microphone as shown in Fig. 1. We use Focusrite Scarlett Solo as the audio interface to amplify the audio signals from the binaural microphone.
@@
TABLE I: Results of the AudioGoal navigation experiment and our model strongly outperforms existing methods.
@@ VI-A Results on Continuous AudioGoal Navigation Benchmark
We first demonstrate the effectiveness of our navigation system on the challenging Continuous AudioGoal navigation benchmark, where the agent moves in a continuous unseen environment to find the location of a ringing telephone sound. For metrics, we use the common Success Rate (SR), success weighted by inverse path length (SPL), and soft SPL. An episode is considered successful when the agent issues the stop action within 1 meter of the goal. SPL [7] is defined as $\mathrm{SPL_{i}}=S_{i}\cdot l_{i}/\mathrm{max}(p_{i},l_{i})$, where $i$ denotes the index of the episode, $S=1$ when the episode is successful and $S=0$ otherwise, $l$ denotes the length of the shortest path between the agent and the audio goal, and $p$ denotes the length of the actual path taken by the agent in the episode. Soft SPL is a variation of SPL where $S_{i}=1$ for all $i$.
@@
Results are shown in Tab. I. Our model strongly outperforms all baselines and ablations. Compared to DDPPO, our model is more efficient due to its hierarchical nature since the DDPPO model often gets stuck with obstacles and corners. Direction Follower and Gan et al. predict the goal direction/location directly, which is however ill-posed when goals are in some other room at a distance from the robot. As a result, their navigation performance is also pretty poor. For the Beamforming baseline, similar to ours, it also predicts the local direction of arrival of the sound, however, since it is not robust to reverberation and noise, it performs quite badly. Lastly, the two ablations perform comparably to baselines but also underperform the full model, showing it is beneficial to predict the full acoustic field and our model uses visual sensors to understand the environment.
@@ VI-B Experiment 2: Acoustic Field Prediction on Real Data
TABLE II: Results for testing on real acoustic field data.

### v2

@@ V Robot Platform
To deploy our sim2real policy on a real robot, we build our audio-visual robot by equipping a HelloRobot with a 3Dio binaural microphone as shown in Fig. 1. We use Focusrite Scarlett Solo as the audio interface to amplify the audio signals from the binaural microphone.
@@
TABLE I: Results of the AudioGoal navigation experiment. Our model strongly outperforms existing methods.
@@ VI-A Results on Continuous AudioGoal Navigation Benchmark
We first demonstrate the effectiveness of our navigation system on the challenging Continuous AudioGoal navigation benchmark [18], where the agent moves in a continuous unseen environment to find the location of a ringing telephone sound. For metrics, we use the common Success Rate (SR), success weighted by inverse path length (SPL), and soft SPL. An episode is considered successful when the agent issues the stop action within 1 meter of the goal. SPL [7] is defined as $\mathrm{SPL_{i}}=S_{i}\cdot l_{i}/\mathrm{max}(p_{i},l_{i})$, where $i$ denotes the index of the episode, $S=1$ when the episode is successful and $S=0$ otherwise, $l$ denotes the length of the shortest path between the agent and the audio goal, and $p$ denotes the length of the actual path taken by the agent in the episode. Soft SPL is a variation of SPL where $S_{i}=1$ for all $i$.
@@
Results are shown in Tab. I. Our model strongly outperforms all baselines and ablations. Compared to DDPPO, our model is more efficient due to its hierarchical nature since the DDPPO model often gets stuck with obstacles and corners. Direction Follower and Gan et al. predict the goal direction/location directly, which is however ill-posed when goals are in some other room at a distance from the robot. As a result, their navigation performance is also pretty poor. For the Beamforming baseline, similar to ours, it also predicts the local direction of arrival of the sound; however, since it is not robust to reverberation and noise, it performs poorly. Lastly, the two ablations perform comparably to baselines but also underperform the full model, showing it is both beneficial to predict the full acoustic field and leverage visual sensors to understand the environment.
@@ VI-B Experiment 2: Acoustic Field Prediction on Real Data
TABLE II: Results for testing on real acoustic field data.
@@ VII Conclusion
While this work represents an exciting first step, it also has some limitations. First, the validation and test data were collected within the same environment, leaving the generalization to novel acoustic environments yet to be explored. Second, we assume the sound sources to be static, which may not hold in all cases, calling for new solutions to address dynamic objects.
Acknowledgements: UT Austin is supported in part by NSF CCRI and IFML NSF AI Institute. KG is paid as a research scientist by Meta.